OpenStudy (anonymous):

Simplify the expression -5z+4n+18z-10n

pooja195 (pooja195):

Combine like termss

pooja195 (pooja195):

-10n+4n=? 18z-5z=?

OpenStudy (anonymous):

Then ?

pooja195 (pooja195):

do the math.

pooja195 (pooja195):

\(\color{blue}{\text{Originally Posted by}}\) @pooja195 -10n+4n=? 18z-5z=? \(\color{blue}{\text{End of Quote}}\) solve

OpenStudy (anonymous):

Thanks